OpenStudy (anonymous):

Coach Johnson is 10 years older than Tim. When Tim started football 3 years ago, Coach Johnson was twice his age. How old is Tim now?

OpenStudy (lgbasallote):

let T = tim's age since coach johnson is 10 years older than Tim, Coach = T + 10 3 years ago...that means Tim's age would be (T-3) and Coach would be (T+10) - 3 it says that 2 years ago coach was twice Tim's age..this means that coach's age three years ago is twice the age of tim three years ago (t+10) - 3 = 2(t-3) t + 10 - 3 = 2t - 6 t+7 = 2t - 6 7+6 = 2t-t understood?

OpenStudy (swag):

hes 13 bro :)